Calculate the total volume of solution when
4 mol of water is mixed with 1 mol of ethanol.
Partial molar volumes of water and ethanol are
1.9 x 10^(-5 ) m3/mol 5.4 x 10^(-5) m3/mol
respectively
1
Expert's answer
2018-01-04T03:38:56-0500
V = n · Vpart
V (H2O) = 4 · 1.9 x 10^-5 = 7.6 x 10-5 m3
V (C2H5OH) = 1 · 5.4 x 10^-5 = 5.4 x 10^-5 m3
Vtotal = 7.6 x 10^-5 + 5.4 x 10^-5 = 0.00013 m3